OpenStudy (anonymous):

I'm having a hard time solving this, I know that this equation is true but I can't work it though to check. 1/8 = 3(7/40) -2/5 Any help would be great thanks. I'm having a hard time solving this, I know that this equation is true but I can't work it though to check. 1/8 = 3(7/40) -2/5 Any help would be great thanks. @Mathematics

OpenStudy (anonymous):

1/8=21/40 -2/5 5/40=21*40 - 16/40 5/40=5/40

OpenStudy (anonymous):

Oh I didn't realise I could increase 1/8 thank you